OpenStudy (anonymous):

Please helpp Given a dilation around the orgin, what is the scale factor K? Do, K= (2,4) → (3,6) a. 2 b. 3.2 c. 2/3 d. 3

OpenStudy (anonymous):

The total linear displacement is the length: \(\Delta l=\sqrt{\Delta x^2 + \Delta y^2} = \sqrt{1^2+2^2}\) The dilation is the displacement divided by the magnitude of the original vector.

OpenStudy (anonymous):

Do you know how to find a vectors magnitude?

OpenStudy (anonymous):

Not really

OpenStudy (anonymous):

You need to study. I think the answer should be 1.5

OpenStudy (anonymous):

Alright. thank you.

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

http://www.mathopenref.com/dilate.html